jhonyy9:

- may be proving that an imaginary number on an imaginary power is real ? for example : i^(1/i) or why not ?

imqwerty:

\(\large{z_1 = re^{i\theta}}\) \(z_2 = a+ ib\) \(\Large {z_1^{z_2}}\) \(\Large {\left(re^{i\theta}\right)^{a+ib}}\) you can simplify it further to find out the conditions for it to be real/imaginary